OpenStudy (anonymous):

Out of 1,280 students. 25% are juniors. Of those 30% either took music or art, how many juniors took either music or art?

OpenStudy (anonymous):

You have to make a venn diagram

OpenStudy (ahsome):

Ok, first find the number of Juniors \[Juniors = Total * 0.25\] This equals to 320 students THEN, find the number of Juniors who did music or art \[Music-and-art = Juniors * 0.3\] This is 96 students

OpenStudy (ahsome):

To convert percentage to decimal, do this\[\frac{ percentage }{ 100 }\] In this instance, \[\frac{ 25}{ 100 }=0.25\] \[\frac{ 30}{ 100}=0.3\]

OpenStudy (anonymous):

omg! I got the correct answer. thank you Ahsome . I was checking to see if my answer was my correct or not.

OpenStudy (ahsome):

@azulapaws No problem bro :)